# Counting integer matrices with square-free determinants

### v1

@@ 5.2. Main term
$\mathscr{M}$ $=\sum_{1\leq d\leq\Delta}\mu(d)\left(N_{n}(d^{2})\frac{H^{n^{2}}}{d^{2n^{2}}}+O\left(d^{2n^{2}-4(n+1)/(n+3)+o(1)}\right)\right)$
$=H^{n^{2}}\sum_{1\leq d\leq\Delta}N_{n}(d^{2})\frac{\mu(d)}{d^{2n^{2}}}+O\left(\Delta^{2n^{2}-(3n+1)/(n+3)+o(1)}\right).$
@@
$\mathscr{M}$ $=H^{n^{2}}\sum_{d=1}^{\infty}N_{n}(d^{2})\frac{\mu(d)}{d^{2n^{2}}}$
@@
$=H^{n^{2}}\sum_{d=1}^{\infty}N_{n}(d^{2})\frac{\mu(d)}{d^{2n^{2}}}+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{2n^{2}-(3n+1)/(n+3)+o(1)}\right).$
@@
(5.2) $\mathscr{M}=\mathfrak{S}_{n}H^{n^{2}}+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{2n^{2}-(3n+1)/(n+3)+o(1)}\right).$
@@ 5.3. Error term
$N_{n}(d^{2};H)\leq\sum_{|a|\leq n!H^{n}/d^{2}}\#\left\{A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right):\leavevmode\nobreak\ \det A=ad^{2}\right\}.$
@@ 5.4. Final optimisation
$S_{n}(H)=\mathfrak{S}_{n}H^{n^{2}}+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{2n^{2}-(3n+1)/(n+3)+o(1)}\right),$
@@ 6.1. Initial split
where, as before, $\mu(d)$ denotes the Möbius function, see, for example, [5, Equation (16.1.3) and Section 16.3]. Therefore, we obtain
@@ 6.2. Main term
$\mathscr{M}$ $=\sum_{1\leq d\leq\Delta}\mu(d)\left(N_{n}(d)\frac{H^{n^{2}}}{d^{n^{2}+1}}+O\left(d^{n^{2}-2+1/n+o(1)}\right)\right)$
$=H^{n^{2}}\sum_{1\leq d\leq\Delta}N_{n}(d)\frac{\mu(d)}{d^{n^{2}+1}}+O\left(\Delta^{n^{2}-1+1/n+o(1)}\right).$
@@
$\mathscr{M}$ $=H^{n^{2}}\sum_{d=1}^{\infty}N_{n}(d)\frac{\mu(d)}{d^{n^{2}+1}}$
@@
$=H^{n^{2}}\sum_{d=1}^{\infty}N_{n}(d)\frac{\mu(d)}{d^{n^{2}+1}}+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{n^{2}-1+1/n+o(1)}\right).$
@@
(6.2) $\mathscr{M}=\sigma_{n}H^{n^{2}}+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{n^{2}-1+1/n+o(1)}\right).$
@@ 6.4. Final optimisation
$\varPhi_{n}(H)=\sigma_{n}H^{n^{2}}+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{n^{2}-1+1/n+o(1)}\right),$

### v2

@@ 5.2. Main term
$\mathscr{M}$ $=\sum_{1\leq d\leq\Delta}\mu(d)\left(N_{n}(d^{2})\frac{(2H+1)^{n^{2}}}{d^{2n^{2}}}+O\left(d^{2n^{2}-4(n+1)/(n+3)+o(1)}\right)\right)$
$=(2H+1)^{n^{2}}\sum_{1\leq d\leq\Delta}N_{n}(d^{2})\frac{\mu(d)}{d^{2n^{2}}}+O\left(\Delta^{2n^{2}-(3n+1)/(n+3)+o(1)}\right).$
@@
$\mathscr{M}$ $=(2H+1)^{n^{2}}\sum_{d=1}^{\infty}N_{n}(d^{2})\frac{\mu(d)}{d^{2n^{2}}}$
@@
$=(2H+1)^{n^{2}}\sum_{d=1}^{\infty}N_{n}(d^{2})\frac{\mu(d)}{d^{2n^{2}}}$
$\qquad\qquad\qquad\qquad+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{2n^{2}-(3n+1)/(n+3)+o(1)}\right).$
@@
(5.2) $\begin{split}\mathscr{M}=\mathfrak{S}_{n}&(2H+1)^{n^{2}}\\ &+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{2n^{2}-(3n+1)/(n+3)+o(1)}\right).\end{split}$
@@ 5.3. Error term
$N_{n}(d^{2};H)\leq\sum_{|a|\leq n!H^{n}/d^{2}}\#\left\{A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right):\penalty 10000\ \det A=ad^{2}\right\}.$
@@ 5.4. Final optimisation
$S_{n}(H)=2^{n^{2}}\mathfrak{S}_{n}H^{n^{2}}+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{2n^{2}-(3n+1)/(n+3)+o(1)}+H^{n^{2}-1}\right),$
@@ 6.1. Initial split
where, as before, $\mu(d)$ denotes the Möbius function, see, for example, [6, Equation (16.1.3) and Section 16.3]. Therefore, we obtain
@@ 6.2. Main term
$\mathscr{M}$ $=\sum_{1\leq d\leq\Delta}\mu(d)\left(N_{n}(d)\frac{(2H+1)^{n^{2}}}{d^{n^{2}+1}}+O\left(d^{n^{2}-2+1/n+o(1)}\right)\right)$
$=(2H+1)^{n^{2}}\sum_{1\leq d\leq\Delta}N_{n}(d)\frac{\mu(d)}{d^{n^{2}+1}}+O\left(\Delta^{n^{2}-1+1/n+o(1)}\right).$
@@
$\mathscr{M}$ $=(2H+1)^{n^{2}}\sum_{d=1}^{\infty}N_{n}(d)\frac{\mu(d)}{d^{n^{2}+1}}$
@@
$=(2H+1)^{n^{2}}\sum_{d=1}^{\infty}N_{n}(d)\frac{\mu(d)}{d^{n^{2}+1}}+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{n^{2}-1+1/n+o(1)}\right).$
@@
(6.2) $\mathscr{M}=\sigma_{n}(2H+1)^{n^{2}}+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{n^{2}-1+1/n+o(1)}\right).$
@@ 6.4. Final optimisation
$\varPhi_{n}(H)=2^{n^{2}}\sigma_{n}H^{n^{2}}+O\left(H^{n^{2}}\Delta^{-1+o(1)}+\Delta^{n^{2}-1+1/n+o(1)}+H^{n^{2}-1}\right),$
@@ Acknowledgement
The authors would like to thank Alan Haynes for pointing out at a recurring oversight in the initial version of this paper.